Circular No. 7863

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bu IN NGC 4242 T. Puckett, Mountain Town, GA; and S. Gauthier, London, ON, re-

port the discovery of an apparent supernova (mag 15.5) on an unfiltered CCD frame (limiting mag 19.2) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Mar. 28.26 UT. The candidate is located at α = 12^{h}17^{m}37^{s}.18, δ = +45^{o}38′47′′.4 (equinox 2000.0), which is 72′′.9 east and 97′′.6 north of the center of NGC 4242. The new object was also present and confirmed on unfiltered CCD frames taken on Mar. 29.02. SN 2002bu does not appear on unfiltered images taken by Puckett on 2001 Feb. 21 and Mar. 14 (limiting mag ∼ 20.5) or on Palomar Sky Survey images taken on 1996 Mar. 17 (limiting mag ∼ 21.0), 1990 Jan. 29 (limiting mag ∼ 21.0), and 1955 Mar. 1 (limiting mag ∼ 20.0).

COMETS C/1998 A4, C/1999 P6, C/1999 P7, C/1999 P8, C/1999 P9

Further to IAUC 7862, D. Hammer reports his measurements of additional comets found by R. Kracht on SOHO website C2 coronagraph images. Reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below. C/1999 P7 dynamically belongs to the Meyer group (cf. IAUC 7832), and the rest belong to the Marsden group.

Comet  1998 UT  α_{2000}  δ_{2000}  MPEC
C/1998 A4 Jan. 10.978  19^{h}26.^{m}5  −20^{o}59′  2002-F70
Comet  1999 UT  α_{2000}  δ_{2000}  MPEC
C/1999 P6 Aug. 5.060  8^{h}52.^{m}9  +17^{o}49′  2002-F43
C/1999 P7  12.996  9 23.3  +16 15  2002-F43
C/1999 P8  15.004  9 30.3  +15 12  2002-F43
C/1999 P9  15.229  9 35.2  +15 15  2002-F43

SUPERNOVA 2002bo IN NGC 3190 K. Sárneczky, Budapest, reports that SN 2002bo (cf. IAUC 7847) is

not present on two 300-s R-filtered CCD images (limiting mag ∼ 19.5) taken by Z. Bebesi and himself on Mar. 4.0 UT with the 0.60-m Schmidt telescope at Konkoly Observatory.

2002 March 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green